# AWM[Uncaptioned image]: Accurate Weight-Matrix Fingerprint for Large Language Models

### v2

@@ 3 Preliminary
Most Large Language Models (LLMs) follows a decoder-only Transformer architecture (radford2018improving). The details of various LLMs may differ, yet the Transformer blocks are similar. In particular, a Transformer block in an LLM usually consists of residual connections (he2016deep), Root Mean Square Normalization (RMSNorm, zhang2019root), the self-attention mechanism (lin2017structured), Rotary Position Embedding (RoPE, su2024roformer), and a feed-forward network (FFN). We denote by $\mathbb{W}_{A}$ the set of an $L$-layered LLM A’s weights, and $\mathbb{W}_{A,\text{partial}}=\{W_{A,\text{emb}}\}\bigcup_{l=1}^{L}\{W_{A,i}^{(l)}\mid i\in\{\text{Q},\text{K}\}\}$ the set of word embeddings and $\text{Q},\text{K}$ matrices, where $W_{A,\text{emb}}$ is the word embeddings and $W_{A,\text{Q}}^{(l)},W_{A,\text{K}}^{(l)}$ are the $\text{Q},\text{K}$ matrices at the $l$-th layer. We defer the rest of the notations to Appendix D.1.
@@
Even with semi-orthogonal $U_{1}$ and $U_{2}$ where $U_{1}\in\mathbb{R}^{n_{1}\times n_{1}^{\prime}}$, $U_{2}\in\mathbb{R}^{n_{2}\times n_{2}^{\prime}}$, $n_{1}>n_{1}^{\prime}$, $n_{2}>n_{2}^{\prime}$, CKA still preserves input similarity to some extent (kang2025spectral; chun2025estimating). Nevertheless, the standard HSIC estimator converges at rate $1/\sqrt{m}$ and shows finite-sample bias (gretton2005measuring; murphy2024correcting). To this end, an unbiased version (song2007supervised) is proposed and extends the value of CKA from $[0,1]$ to [1,1][-1,1]. Additionally, linear kernels are often selected in CKA due to their computational efficiency and similar performance to other kernels (kornblith2019similarity). We give the formal definition of CKA in Appendix D.2, and the proof of Theorem 3.1 in Appendix E.1.
@@ 4.2 Residual Connections: Passing Manipulations Forward
Proposition 4.2 (Proof in Appendix E.2).
@@
Compared to zeng2024huref, Section 4.2 shows that residual connections are more vulnerable under a component-wise view, since constituent functions can propagate manipulations in various ways. However, nonlinearities of the functions, especially those within the self-attention mechanism, pose constraints for manipulations on both inputs and weights. We next split the self-attention mechanism into two constituent functions, RMSNorm and attention (see Section D.1), and show how these effects arise and are constrained these two functions with a focus on RMSNorm, RoPE and attention scores.
@@ 4.3 RMSNorm: A Constraint on Embedding Manipulations
Let models $A$ and $B$ share the same [^1] The conclusions generalize to models with different number of layers (possibly a result of pruning). See Appendix E.3architecture. Let $c\neq 0$ be a scalar, $P$ be a (partial) permutation matrix, and $D$ be a signature matrix. Then, $R_{B,\text{emb}}=cPD$ can be recovered after RMSNorm in model B if related RMSNorm parameters are adjusted.
Theorem 4.3 indicates that RMSNorm is susceptible to embedding manipulations composed of constant multiplications, permutations and sign flips. On the other hand, other manipulations on word embeddings are generally neither commutative with RMSNorm nor invariant to it, which potentially brings a constraint to manipulations. We provide a proof for Theorem 4.3 in Appendix E.3, along with a discussion on other manipulations on word embeddings.
@@ 4.4 RoPE and Attention Scores: Boundaries for Q/K Manipulations
A proof for Theorem 4.4 is provided in Appendix E.4, where we also show how manipulations on inputs are recovered after V, O matrices to satisfy Section 4.2. These manipulations preserve the attention score values of model A in the suspicious model B. However, they can greatly change the weights of the original model A, bringing difficulty to the development of detection methods.
@@ 4.5 Potential Attacks
Here $E$ collects post-training, continual pre-training, pruning, upcycling, multimodal adaptation, or related adjustments. The nonlinear usage of $\mathbb{W}_{A,\text{partial}}$, especially through the $\text{Q},\text{K}$ matrices, substantially limits manipulation complexity. Consequently, checking similarity between $\mathbb{W}_{A,\text{partial}}$ and $\mathbb{W}_{B,\text{partial}}$ is typically sufficient for detection. The remaining avenues targeting $\mathbb{W_{A}}/\mathbb{W}_{A,\text{partial}}$ are deferred to Appendix E.5, where we also discuss how such transformations can recover Transformer block outputs in light of Section 4.2.
@@ 5 Methodology
Despite accounting for permutation and signature manipulations, the orthogonal transformations $U_{B,i}^{(l)}$ remain challenging. They introduce a substantial nuisance parameter burden: if $W_{A,i}^{(l)}\in\mathbb{R}^{d\times n}$, then orthogonal $U_{B,i}^{(l)}\in\mathbb{R}^{d\times d}$ contributes $d^{2}$ parameters, which is prohibitive when $d\approx n$ and can undermine robustness by adding parameters to detections (simmons2011false). We therefore use Central Kernel Alignment (CKA) as a parameter-free similarity metric: by Theorem 3.1, CKA is invariant to orthogonal transforms and constant rescaling. While this invariance does not extend to semi-orthogonal $U_{B,i}^{(l)}$ (e.g., induced by pruning), CKA remains effective to a meaningful extent in such cases, relieving the need to explicitly reconstruct $U_{B,i}^{(l)}$. To further mitigate biases (gretton2005measuring; murphy2024correcting), we adopt the unbiased variant (song2007supervised), termed UCKA (see Appendix D.2).
@@ 6.3 Robustness Verification on 60 Offspring Models
Table 2: Absolute Z-scores ($\uparrow$) of positive samples for HuRef, REEF, and our method. The full names of abbreviations and related model details are provided in Appendix G.
@@ 7 Conclusions
Appendix A Ethics Statement
The primary motivation for this work is to protect the intellectual property of large language models (LLMs), thereby promoting fairness and accountability within the AI community. We acknowledge the potential societal impact of such a technology, particularly the risk of false accusations of model theft. Therefore, a core design principle of our method is to achieve an exceptionally low false-positive rate, as empirically demonstrated in our experiments. By establishing a high-fidelity verification system, we aim to foster a more transparent and trustworthy open-source ecosystem. All models used in this study are publicly available, and our research does not involve any private or sensitive user data.
Appendix B Reproducibility Statement
To ensure the reproducibility of our results, we have included the complete source code in the supplementary material. We also provided a detailed description of our methodology in Section 5, including a step-by-step algorithm. All models used for evaluation are publicly available from sources such as the Hugging Face Hub, and a comprehensive list mapping our model abbreviations to their full names is included in Appendix G. A public code repository will be made available upon publication.
Appendix C The Use of Large Language Models
The core research and analysis presented in this manuscript were conducted without the use of Large Language Models (LLMs). An LLM was utilized exclusively to improve the language and clarity of the text.
Appendix D Definitions and Notations
D.1 LLM Architecture
@@
Definition D.2 (Self-Attention with RoPE).
@@
RoPE’s definition in Section D.1 is slightly different from their implementations, but still an equivalent version to su2024roformer.
Definition D.3 (Feed-forward Networks).
@@
We abuse the notations in Section D.1 by denoting $\text{RMSNorm}_{\text{attn}}^{(l)}(X)\triangleq\text{RMSNorm}(X,\omega_{A,\text{attn}}^{(l)},\epsilon_{A})$ as the attention norm and $\text{RMSNorm}_{\text{ffn}}^{(l)}(X)\triangleq\text{RMSNorm}(X,\omega_{A,\text{ffn}}^{(l)},\epsilon_{A})$ the ffn norm at layer $l$ of model A. Then, the definition of a Transformer block at layer $l$ of model A can be summartized as follows.
Definition D.4 (Transformer Block).
@@
D.2 Central Kernel Alignment (CKA)
@@
Appendix E Proofs and Discussions
E.1 Proof for Theorem 3.1
We give the proof based on the definition of CKA with linear kernels in Appendix D.2. Given input matrices $X_{1},X_{2}$, orthogonal matrices $U_{1},U_{2}$ and constants $c_{1},c_{2}$, the kernel functions yield
@@ Appendix E Proofs and Discussions
E.3 Proof for Theorem 4.3 and Discussions on Other Input Manipulations
We first show a proof for the claims in Theorem 4.3. The core lies in preserving the diagonal structure of RMSNorm parameters in Section D.1. Given the input $X$ to layer $l$, the manipulations change it to $cXPD$ and fed the manipulated input into the attention norm of model B. In order to recover the manipulations on inputs after RMSNorm, i.e. for any manipulation $T$ in the input $X$ (an example is $T(X)=cXPD$),
@@
E.4 Proof for Theorem 4.4 and Related Discussions
We first provide the proof for Theorem 4.4 based on Section D.1.
@@
E.5 How Potential Attacks in Section 4.5 Recovers the Output of Transformers
We provide an illustration based on Section D.1, Section D.1 and Section D.1. Previous parts have demonstrated how the manipulations on inputs pass through self-attention. Hence, it suffices to show how the manipulations pass through FFN and yield an output identical to model A after the language model head.
@@
By Section D.1 and Section D.1, the manipulation is propagated through transformer layers. At the language model head, we denote $W_{B,\text{lm}}=c^{-1}W_{A,\text{lm}}PD$ and abuse $X^{\prime}=cXPD$ as the manipulated input. Then,
@@
Appendix F Ablation Studies
F.1 Number of Overlapping Vocabulary Tokens
@@ F.2 CKA Ablations
Appendix G Model Details
@@
Appendix H Implementation Details
@@
Appendix I Empirical Validation of Robustness against Weight Manipulations

### v3

@@ 3 Preliminary
Most Large Language Models (LLMs) follows a decoder-only Transformer architecture (radford2018improving). The details of various LLMs may differ, yet the Transformer blocks are similar. In particular, a Transformer block in an LLM usually consists of residual connections (he2016deep), Root Mean Square Normalization (RMSNorm, zhang2019root), the self-attention mechanism (lin2017structured), Rotary Position Embedding (RoPE, su2024roformer), and a feed-forward network (FFN). We denote by $\mathbb{W}_{A}$ the set of an $L$-layered LLM A’s weights, and $\mathbb{W}_{A,\text{partial}}=\{W_{A,\text{emb}}\}\bigcup_{l=1}^{L}\{W_{A,i}^{(l)}\mid i\in\{\text{Q},\text{K}\}\}$ the set of word embeddings and $\text{Q},\text{K}$ matrices, where $W_{A,\text{emb}}$ is the word embeddings and $W_{A,\text{Q}}^{(l)},W_{A,\text{K}}^{(l)}$ are the $\text{Q},\text{K}$ matrices at the $l$-th layer. We defer the rest of the notations to Appendix A.1.
@@
Even with semi-orthogonal $U_{1}$ and $U_{2}$ where $U_{1}\in\mathbb{R}^{n_{1}\times n_{1}^{\prime}}$, $U_{2}\in\mathbb{R}^{n_{2}\times n_{2}^{\prime}}$, $n_{1}>n_{1}^{\prime}$, $n_{2}>n_{2}^{\prime}$, CKA still preserves input similarity to some extent (kang2025spectral; chun2025estimating). Nevertheless, the standard HSIC estimator converges at rate $1/\sqrt{m}$ and shows finite-sample bias (gretton2005measuring; murphy2024correcting). To this end, an unbiased version (song2007supervised) is proposed and extends the value of CKA from $[0,1]$ to [1,1][-1,1]. Additionally, linear kernels are often selected in CKA due to their computational efficiency and similar performance to other kernels (kornblith2019similarity). We give the formal definition of CKA in Appendix A.2, and the proof of Theorem 3.1 in Appendix B.1.
@@ 4.2 Residual Connections: Passing Manipulations Forward
Proposition 4.2 (Proof in Appendix B.2).
@@
Compared to zeng2024huref, Section 4.2 shows that residual connections are more vulnerable under a component-wise view, since constituent functions can propagate manipulations in various ways. However, nonlinearities of the functions, especially those within the self-attention mechanism, pose constraints for manipulations on both inputs and weights. We next split the self-attention mechanism into two constituent functions, RMSNorm and attention (see Section A.1), and show how these effects arise and are constrained these two functions with a focus on RMSNorm, RoPE and attention scores.
@@ 4.3 RMSNorm: A Constraint on Embedding Manipulations
Let models $A$ and $B$ share the same [^1] The conclusions generalize to models with different number of layers (possibly a result of pruning). See Appendix B.3architecture. Let $c\neq 0$ be a scalar, $P$ be a (partial) permutation matrix, and $D$ be a signature matrix. Then, $R_{B,\text{emb}}=cPD$ can be recovered after RMSNorm in model B if related RMSNorm parameters are adjusted.
Theorem 4.3 indicates that RMSNorm is susceptible to embedding manipulations composed of constant multiplications, permutations and sign flips. On the other hand, other manipulations on word embeddings are generally neither commutative with RMSNorm nor invariant to it, which potentially brings a constraint to manipulations. We provide a proof for Theorem 4.3 in Appendix B.3, along with a discussion on other manipulations on word embeddings.
@@ 4.4 RoPE and Attention Scores: Boundaries for Q/K Manipulations
A proof for Theorem 4.4 is provided in Appendix B.4, where we also show how manipulations on inputs are recovered after V, O matrices to satisfy Section 4.2. These manipulations preserve the attention score values of model A in the suspicious model B. However, they can greatly change the weights of the original model A, bringing difficulty to the development of detection methods.
@@ 4.5 Potential Attacks
Here $E$ collects post-training, continual pre-training, pruning, upcycling, multimodal adaptation, or related adjustments. The nonlinear usage of $\mathbb{W}_{A,\text{partial}}$, especially through the $\text{Q},\text{K}$ matrices, substantially limits manipulation complexity. Consequently, checking similarity between $\mathbb{W}_{A,\text{partial}}$ and $\mathbb{W}_{B,\text{partial}}$ is typically sufficient for detection. The remaining avenues targeting $\mathbb{W_{A}}/\mathbb{W}_{A,\text{partial}}$ are deferred to Appendix B.5, where we also discuss how such transformations can recover Transformer block outputs in light of Section 4.2.
@@ 5 Methodology
Despite accounting for permutation and signature manipulations, the orthogonal transformations $U_{B,i}^{(l)}$ remain challenging. They introduce a substantial nuisance parameter burden: if $W_{A,i}^{(l)}\in\mathbb{R}^{d\times n}$, then orthogonal $U_{B,i}^{(l)}\in\mathbb{R}^{d\times d}$ contributes $d^{2}$ parameters, which is prohibitive when $d\approx n$ and can undermine robustness by adding parameters to detections (simmons2011false). We therefore use Central Kernel Alignment (CKA) as a parameter-free similarity metric: by Theorem 3.1, CKA is invariant to orthogonal transforms and constant rescaling. While this invariance does not extend to semi-orthogonal $U_{B,i}^{(l)}$ (e.g., induced by pruning), CKA remains effective to a meaningful extent in such cases, relieving the need to explicitly reconstruct $U_{B,i}^{(l)}$. To further mitigate biases (gretton2005measuring; murphy2024correcting), we adopt the unbiased variant (song2007supervised), termed UCKA (see Appendix A.2).
@@ 6.3 Robustness Verification on 60 Offspring Models
Table 2: Absolute Z-scores ($\uparrow$) of positive samples for HuRef, REEF, and our method. The full names of abbreviations and related model details are provided in Appendix D.
@@ 7 Conclusions
8 Acknowledgements
This work is sponsored by the National Key Research and Development Program of China (No. 2023ZD0121402) and the National Natural Science Foundation of China (NSFC) grant (No. 62576211).
Appendix A Definitions and Notations
A.1 LLM Architecture
@@
Definition A.2 (Self-Attention with RoPE).
@@
RoPE’s definition in Section A.1 is slightly different from their implementations, but still an equivalent version to su2024roformer.
Definition A.3 (Feed-forward Networks).
@@
We abuse the notations in Section A.1 by denoting $\text{RMSNorm}_{\text{attn}}^{(l)}(X)\triangleq\text{RMSNorm}(X,\omega_{A,\text{attn}}^{(l)},\epsilon_{A})$ as the attention norm and $\text{RMSNorm}_{\text{ffn}}^{(l)}(X)\triangleq\text{RMSNorm}(X,\omega_{A,\text{ffn}}^{(l)},\epsilon_{A})$ the ffn norm at layer $l$ of model A. Then, the definition of a Transformer block at layer $l$ of model A can be summartized as follows.
Definition A.4 (Transformer Block).
@@
A.2 Central Kernel Alignment (CKA)
@@
Appendix B Proofs and Discussions
B.1 Proof for Theorem 3.1
We give the proof based on the definition of CKA with linear kernels in Appendix A.2. Given input matrices $X_{1},X_{2}$, orthogonal matrices $U_{1},U_{2}$ and constants $c_{1},c_{2}$, the kernel functions yield
@@ Appendix B Proofs and Discussions
B.3 Proof for Theorem 4.3 and Discussions on Other Input Manipulations
We first show a proof for the claims in Theorem 4.3. The core lies in preserving the diagonal structure of RMSNorm parameters in Section A.1. Given the input $X$ to layer $l$, the manipulations change it to $cXPD$ and fed the manipulated input into the attention norm of model B. In order to recover the manipulations on inputs after RMSNorm, i.e. for any manipulation $T$ in the input $X$ (an example is $T(X)=cXPD$),
@@
B.4 Proof for Theorem 4.4 and Related Discussions
We first provide the proof for Theorem 4.4 based on Section A.1.
@@
B.5 How Potential Attacks in Section 4.5 Recovers the Output of Transformers
We provide an illustration based on Section A.1, Section A.1 and Section A.1. Previous parts have demonstrated how the manipulations on inputs pass through self-attention. Hence, it suffices to show how the manipulations pass through FFN and yield an output identical to model A after the language model head.
@@
By Section A.1 and Section A.1, the manipulation is propagated through transformer layers. At the language model head, we denote $W_{B,\text{lm}}=c^{-1}W_{A,\text{lm}}PD$ and abuse $X^{\prime}=cXPD$ as the manipulated input. Then,
@@
Appendix C Ablation Studies
C.1 Number of Overlapping Vocabulary Tokens
@@ C.2 CKA Ablations
Appendix D Model Details
@@
Appendix E Implementation Details
@@
Appendix F Empirical Validation of Robustness against Weight Manipulations